# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ B.1 The Case $L=2$
where $\bm{e}_{i}\in\mathbb{R}^{d_{\min}}$ is a standard basis vector with the $i$-th entry being 1 and all other entries being 00. According to 3, one can verify that $\bm{W}(0)\in\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$. Using $y_{i}=\sqrt{\lambda}$, we obtain $\sigma^{*}_{i}\in\{x:x^{3}-\sqrt{\lambda}y_{i}x+\lambda x=0,\ x\geq 0\}=\{0\}$. Then, we compute
@@
where the first inequality follows from Mirsky’s inequality (see Lemma 11). When $t\leq{\delta_{\sigma}}/{3}$, we have
$\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})\leq\|\bm{W}(t)-\bm{W}(0)\|_{F}\leq\frac{\sqrt{2}\delta_{\sigma}}{3}.$
Using this inequality and 2(ii), we have for each $\bar{\bm{\sigma}}^{*}\in\mathcal{A}$ satisfying $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}\neq\widehat{\mathcal{W}}_{\bar{\bm{\sigma}}^{*}}$,
@@
Consequently, we obtain $\mathrm{dist}(\bm{W}(t),\mathcal{W}_{G})=\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})$. This, together with (184) and (185), implies
$\|\nabla G(\bm{W}(t))\|_{F}=2\sqrt{2}|t|^{3}\leq 2^{3/4}\mathrm{dist}^{\frac{3}{2}}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=2^{3/4}\mathrm{dist}^{\frac{3}{2}}(\bm{W}(t),\mathcal{W}_{G}).$
Obviously, the error bound does not hold when $t\to 0$. ∎
B.2 The Case $L\geq 3$
To begin, we show the following lemma to derive the equivalent condition of (76):
Lemma 7.
When $L\geq 3$, (76) holds if and only if for all $0\neq\sigma^{*}\in\mathcal{Y}$, we have $\varphi^{\prime}(\sigma^{*})\neq 0$, where $\cal Y$ and $\varphi(\cdot)$ are defined in (47) and (82), respectively.
@@ B.2 The Case $L\geq 3$
Note that $\varphi^{\prime}(\sigma^{*})=0$ is equivalent to
$\varphi^{\prime}(\sigma^{*})=\frac{L}{\sqrt{\lambda}}\sigma^{*(L-1)}+\sqrt{\lambda}(2-L)\sigma^{*(1-L)}=0.$ (186)
According to $\sigma^{*}\in\mathcal{Y}$, there exists $j\in[d_{\min}]$ such that
$\varphi(\sigma^{*})=\frac{\sigma^{*(2L-1)}+\lambda\sigma^{*}}{\sqrt{\lambda}\sigma^{*(L-1)}}=y_{j}.$ (187)
Combining (186) and (187), we obtain
$y_{j}=\left[\left(\frac{L-2}{L}\right)^{\frac{L}{2(L-1)}}+\left(\frac{L}{L-2}\right)^{\frac{L-2}{2(L-1)}}\right]\lambda^{\frac{1}{2(L-1)}}.$
@@
Conversely, suppose that (76) does not hold. This implies that there exists $j\in[d_{\min}]$ such that $y_{j}=\left[\left(\frac{L-2}{L}\right)^{\frac{L}{2(L-1)}}+\left(\frac{L}{L-2}\right)^{\frac{L-2}{2(L-1)}}\right]\lambda^{\frac{1}{2(L-1)}}$. Using this, one can verify that
$x^{2L-1}-\sqrt{\lambda}y_{j}x^{L-1}+\lambda x=0$ (188)
have a positive root such that
$x^{*}=\left(\frac{\lambda(L-1)}{L}\right)^{\frac{1}{2(L-1)}}\ \text{and}\ \varphi^{\prime}(x^{*})=0.$
By the definition of $\mathcal{Y}$, we have $x^{*}\in\mathcal{Y}$. Then, we complete the proof. ∎
@@
When $L=2$, we compute
@@
Therefore, $\varphi^{\prime}(\sigma^{*})\neq 0$ holds trivially.
Lemma 8.
Suppose that $L\geq 3$ and (76) does not hold. There exists $\bm{\sigma}^{*}\in\cal A$ such that the error bound for the critical point set $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$ of Problem (13) fails to hold, i.e., for any $\kappa,\ \delta>0$, there exist $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\widehat{\mathcal{W}}_{\sigma^{*}})\leq\delta$ such that $\|\nabla G(\bm{W})\|_{F}\leq\kappa\cdot\mathrm{dist}(\bm{W},\mathcal{W}_{G})$.
@@
According to Lemma 7 and the fact that (76) does not hold, there exists $\bm{\sigma}^{*}\in\cal A$ such that $\varphi^{\prime}(\sigma_{i}^{*})=0$ for some $i\in[d_{\min}]$. This implies
$f(\sigma^{*}_{i}):=(\sigma^{*}_{i})^{2L-1}-\sqrt{\lambda}y_{i}(\sigma^{*}_{i})^{L-1}+\lambda\sigma^{*}_{i}=0,\ \varphi(\sigma^{*}_{i})=y_{i}.$
Note that $f(x)=\sqrt{\lambda}x^{L-1}\varphi(x)-\sqrt{\lambda}y_{i}x^{L-1}$ and we compute
$f^{\prime}(\sigma_{i}^{*})=\sqrt{\lambda}(\sigma_{i}^{*})^{L-1}\varphi^{\prime}(\sigma_{i}^{*})=0.$
Now, we define $\bm{W}(t)=(\bm{W}_{1}(t),\cdots,\bm{W}_{L}(t))$ as follow:
$\begin{cases}&\bm{W}_{1}(t)=\bm{Q}_{2}\bm{\Sigma}_{1}(t)\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ &\bm{W}_{l}(t)=\bm{Q}_{l+1}\bm{\Sigma}_{l}(t)\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\bm{Q}_{l}\in\mathcal{O}^{d_{l-1}},\ l=2,\dots,L,\\ &\bm{W}_{L}(t)=\mathrm{BlkDiag}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{L}(t)\bm{Q}_{L}^{T},\\ &\bm{\Sigma}_{l}(t)=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\sigma}^{*}+t\bm{e}_{i}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\ \forall l\in[L],\\ &\bm{O}_{i}\in\mathcal{O}^{h_{i}},\forall i\in[p_{Y}],\ \bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}.\end{cases}$
It follows from 3 that $\bm{W}(0)\in\mathcal{W}_{G}$. Therefore, we obtain for all $l\in[L]$,

### v3

@@ A.1 The Case $L=2$
According to 1, one can verify that $\bm{W}(0)\in\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$. Using $y_{i}=\sqrt{\lambda}$, we obtain
$\sigma^{*}_{i}\in\{x:x^{3}-\sqrt{\lambda}y_{i}x+\lambda x=0,\ x\geq 0\}=\{0\}.$
Then, we compute
@@
where the first inequality uses Mirsky’s inequality (see Lemma D.8). When $t\leq{\delta_{\sigma}}/{3}$, we have $\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})\leq\|\bm{W}(t)-\bm{W}(0)\|_{F}\leq{\sqrt{2}\delta_{\sigma}}/{3}.$ Using this and 1(ii), we have for each $\bar{\bm{\sigma}}^{*}\in\mathcal{A}$ satisfying $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}\neq\widehat{\mathcal{W}}_{\bar{\bm{\sigma}}^{*}}$,
@@
Consequently, we obtain $\mathrm{dist}(\bm{W}(t),\mathcal{W}_{G})=\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})$. This, together with (127) and (128), implies $\|\nabla G(\bm{W}(t))\|_{F}=2\sqrt{2}|t|^{3}\leq 2^{3/4}\mathrm{dist}^{\frac{3}{2}}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=2^{3/4}\mathrm{dist}^{\frac{3}{2}}(\bm{W}(t),\mathcal{W}_{G}).$ Obviously, the error bound does not hold when $t\to 0$. ∎
A.2 The Case $L\geq 3$
We first present the following lemma to derive the equivalent condition to (8).
Lemma A.2.
When $L\geq 3$, (8) holds if and only if $\varphi^{\prime}(\sigma^{*})\neq 0$ for all $\sigma^{*}\in\mathcal{Y}\setminus\{0\}$, where $\cal Y$ and $\varphi(\cdot)$ are defined in (28) and (85), respectively. When $L=2$, we have $\varphi^{\prime}(\sigma^{*})\neq 0$.
@@ A.2 The Case $L\geq 3$
First, consider the case where $L\geq 3$. Suppose that there exists $\sigma^{*}\in\mathcal{Y}\setminus\{0\}$ such that $\varphi^{\prime}(\sigma^{*})=0$. Note that $\varphi^{\prime}(\sigma^{*})=0$ is equivalent to $\varphi^{\prime}(\sigma^{*})=L\sigma^{*(L-1)}/\sqrt{\lambda}+\sqrt{\lambda}(2-L)\sigma^{*(1-L)}=0.$ Using $\sigma^{*}\in\mathcal{Y}$, there exists $j\in[d_{\min}]$ such that $\sigma^{*(2L-1)}+\lambda\sigma^{*}=\sqrt{\lambda}\sigma^{*(L-1)}y_{j}$. Combining the above two equations, we obtain
$y_{j}=\left(\left(\frac{L-2}{L}\right)^{\frac{L}{2(L-1)}}+\left(\frac{L}{L-2}\right)^{\frac{L-2}{2(L-1)}}\right)\lambda^{\frac{1}{2(L-1)}}.$ (129)
@@
Conversely, suppose that (8) does not hold. This implies that there exists $j\in[d_{\min}]$ such that (129) holds. Then, one can verify that $x^{2L-1}-\sqrt{\lambda}y_{j}x^{L-1}+\lambda x=0$ has a positive root $x^{*}=\left({\lambda(L-1)}/{L}\right)^{\frac{1}{2(L-1)}}$ and $\varphi^{\prime}(x^{*})=0$. Using the definition of $\mathcal{Y}$, we have $x^{*}\in\mathcal{Y}$.
When $L=2$, we compute $\varphi^{\prime}(\sigma^{*})=2\sigma^{*}/\sqrt{\lambda}>0$ for all $\sigma^{*}\in\mathcal{Y}\setminus\{0\}.$ So, $\varphi^{\prime}(\sigma^{*})\neq 0$ holds trivially. ∎
Lemma A.3.
Suppose that $L\geq 3$ and (8) does not hold. There exists $\bm{\sigma}^{*}\in\cal A$ such that the error bound for the critical point set $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$ of Problem (12) fails to hold.
Proof.
According to Lemma A.2 and the fact that (8) does not hold, there exists $\bm{\sigma}^{*}\in\cal A$ such that $\varphi^{\prime}(\sigma_{i}^{*})=0$ for some $i\in[d_{\min}]$. This implies $f(\sigma^{*}_{i}):=(\sigma^{*}_{i})^{2L-1}-\sqrt{\lambda}y_{i}(\sigma^{*}_{i})^{L-1}+\lambda\sigma^{*}_{i}=0$ and $\varphi(\sigma^{*}_{i})=y_{i}$. Note that $f(x)=\sqrt{\lambda}x^{L-1}\varphi(x)-\sqrt{\lambda}y_{i}x^{L-1}$ and we compute $f^{\prime}(\sigma_{i}^{*})=\sqrt{\lambda}(\sigma_{i}^{*})^{L-1}\varphi^{\prime}(\sigma_{i}^{*})=0$. Now, we define $\bm{W}(t)=(\bm{W}_{1}(t),\cdots,\bm{W}_{L}(t))$ as follow:
$\begin{cases}&\bm{W}_{1}(t)=\bm{Q}_{2}\bm{\Sigma}_{1}(t)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ &\bm{W}_{l}(t)=\bm{Q}_{l+1}\bm{\Sigma}_{l}(t)\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\\ &\bm{W}_{L}(t)=\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{L}(t)\bm{Q}_{L}^{T},\\ &\bm{\Sigma}_{l}(t)=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma}^{*}+t\bm{e}_{i}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\ \forall l\in[L],\\ &\bm{O}_{i}\in\mathcal{O}^{h_{i}},\forall i\in[p_{Y}],\ \bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}},\ \bm{Q}_{l}\in\mathcal{O}^{d_{l-1}},\ l=2,\dots,L.\end{cases}$
It follows from 1 that $\bm{W}(0)\in\mathcal{W}_{G}$. Therefore, we obtain for all $l\in[L]$,
[…]
Applying the Taylor expansion to $f(\sigma^{*}_{i}+t)$ at $\sigma^{*}$, together with $f(\sigma^{*}_{i})=0$ and $f^{\prime}(\sigma^{*}_{i})=0$, yields that when $t\to 0$, $\|\nabla G(\bm{W}(t))\|_{F}=O(t^{2}).$ We also note that $\mathrm{dist}^{2}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=\|\bm{W}_{1}(t)-\bm{W}_{1}^{*}\|_{F}^{2}+\cdots+\|\bm{W}_{L}(t)-\bm{W}_{L}^{*}\|_{F}^{2}\geq Lt^{2},$ where the inequality follows from Weyl’s inequality. Using the same argument in Lemma A.1, we conclude that $\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=\mathrm{dist}(\bm{W}(t),\mathcal{W}_{G})$ when $t$ is sufficiently small. Then we have $\left\|\nabla G\left(\bm{W}(t)\right)\right\|_{F}=O\left(\mathrm{dist}^{2}(\bm{W}(t),\mathcal{W}_{G})\right)$, which implies that the error bound fails to hold. ∎
@@
Under Assumption 1, it follows from 2 that Assumption 2 is a sufficient condition for the error bound to hold for the critical point set $\mathcal{W}_{G}$. According to Lemma A.1 and Lemma A.3, we conclude that Assumption 2 is also a necessary condition. Therefore, we establish the necessary and sufficient condition unde which the error bound holds.
Appendix B Supplementary Results and Proofs for Section 3
B.1 Proof of 1
Proof.
(i) Obviously, the “only if” direction is trivial since $\bm{\sigma}\in\mathcal{W}_{\bm{\sigma},\bm{\Pi}}$ and $\bm{\sigma}^{\prime}\in\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$ share the same non-increasing singular values. It remains to prove the “if” direction. Suppose that $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$ and $(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})\in\mathcal{B}$ satisfy $\bm{\sigma}=\bm{\sigma}^{\prime}$. There exist $\bm{a},\bm{a}^{\prime}\in\mathcal{A}$ such that
This, together with $\bm{\sigma}=\bm{\sigma}^{\prime}$, implies that $\bm{a},\bm{a}^{\prime}$ have the same positive elements but in a different order. Consider the following equation from $\mathcal{A}$: $x_{i}^{2L-1}-\sqrt{\lambda}y_{i}x_{i}^{L-1}+\lambda x_{i}=0,\ x_{i}\geq 0,\ \forall i\in[d_{\min}].$ Note that if $x_{i}=x_{j}>0$, we have $y_{i}=y_{j}$, which implies that if $y_{i}\neq y_{j}$, we have $x_{i}\neq x_{j}$ when $x_{i},x_{j}>0$. Since $\bm{\sigma}=\bm{\sigma}^{\prime}$ and each element of $\bm{a}$ and $\bm{a}^{\prime}$ is obtained by solving the above equation, along with the fact that different $y_{i}$’s correspond to equations with no common positive root and the partition of $(y_{1},\ldots,y_{d_{\min}})$ in (4), yields that the elements of $\bm{a}$ and $\bm{a}^{\prime}$ in each partition of the form $\bm{Y}$ differ only in their order. Therefore,there exist $\bm{P}_{i}\in\mathcal{P}^{h_{i}}$ for all $i\in[p_{Y}]$ and $\bm{P}_{p_{Y}+1}\in\mathcal{P}^{d_{\min}-r_{Y}}$ such that $\mathrm{diag}(\bm{a})=\mathrm{BlkDiag}(\bm{P}_{1},\ldots,\bm{P}_{p_{Y}},\bm{P}_{p_{Y}+1})\mathrm{diag}(\bm{a}^{\prime})\mathrm{BlkDiag}(\bm{P}_{1}^{T},\ldots,\bm{P}_{p_{Y}}^{T},\bm{P}_{p_{Y}+1}^{T}).$ Substituting (130) into the above equality, together with $\bm{\sigma}=\bm{\sigma}^{\prime}$, yields
$\bm{\Pi}^{T}\mathrm{diag}(\bm{\sigma})\bm{\Pi}=\mathrm{BlkDiag}(\bm{P}_{1},\ldots,\bm{P}_{p_{Y}},\bm{P}_{p_{Y}+1})\bm{\Pi}^{\prime T}\mathrm{diag}(\bm{\sigma})\bm{\Pi}^{\prime}\mathrm{BlkDiag}(\bm{P}_{1}^{T},\ldots,\bm{P}_{p_{Y}}^{T},\bm{P}_{p_{Y}+1}^{T}).$ (131)
$\tilde{\bm{Q}}_{l}:=\bm{Q}_{l}\mathrm{BlkDiag}(\bm{\Pi},\bm{I})\mathrm{BlkDiag}\left(\bm{P}_{1},\ldots,\bm{P}_{p_{Y}},\bm{P}_{p_{Y}+1},\bm{I}\right)\mathrm{BlkDiag}(\bm{\Pi}^{{}^{\prime}T},\bm{I})\in\mathcal{O}^{d_{l-1}},\ \forall l\in\{2,\dots,L\}.$ (132)
Moreover, we have $\bm{\Sigma}_{l}\mathrm{BlkDiag}(\bm{\Pi},\bm{I})=\mathrm{BlkDiag}(\mathrm{diag}(\bm{\sigma})\bm{\Pi},\bm{0}),\ \forall l\in[L].$ Then, we have
$\textstyle\bm{W}_{1}$ $=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{\Pi},\bm{I}\right)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right)$
$=\tilde{\bm{Q}}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{\Pi}^{\prime},\bm{I}\right)\mathrm{BlkDiag}(\bm{P}_{1}^{T}\bm{O}_{1},\ldots,\bm{P}_{p_{Y}}^{T}\bm{O}_{p_{Y}},\mathrm{BlkDiag}(\bm{P}_{p_{Y}+1}^{T},\bm{I}_{d_{0}-d_{\min}})\bm{O}_{p_{Y}+1}),$
$\bm{W}_{L}=\mathrm{BlkD}\left(\bm{O}_{1}^{T}\bm{P}_{1},\dots,\bm{O}_{p_{Y}}^{T}\bm{P}_{Y},\widehat{\bm{O}}_{p_{Y}+1}^{T}\mathrm{BlkDiag}(\bm{P}_{p_{Y}+1},\bm{I})\right)\mathrm{BlkD}\left(\bm{\Pi}^{\prime T},\bm{I}\right)\bm{\Sigma}_{L}\tilde{\bm{Q}}_{L}^{T},$
Therefore, we obtain that $(\bm{W}_{1},\ldots,\bm{W}_{L})\in\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$ and thus $\mathcal{W}_{\bm{\sigma},\bm{\Pi}}\subseteq\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$. Applying the same argument, we also have $\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}\subseteq\mathcal{W}_{\bm{\sigma},\bm{\Pi}}$. Therefore, we have $\mathcal{W}_{\bm{\sigma},\bm{\Pi}}=\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$.
@@ B.1 Proof of 1
B.2 Auxiliary Lemma for Proving Lemma 3
Lemma B.4.
Let $a>0$ be a constant, $\bm{\Sigma}\in\mathbb{R}^{n\times n}$ be a diagonal matrix, and $\bm{Q}\in\mathcal{O}^{n}$ be an orthogonal matrix. Then, if $\|\bm{\Sigma}-a\bm{I}\|\leq a/2$, we have $\|\bm{Q}\bm{\Sigma}^{2}-\bm{\Sigma}^{2}\bm{Q}\|_{F}\geq a\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\|_{F}.$
@@ B.2 Auxiliary Lemma for Proving Lemma 3
For ease of exposition, let $\bm{\Delta}:=\bm{\Sigma}-a\bm{I}$. We compute
$\textstyle\|\bm{Q}\bm{\Sigma}^{2}-\bm{\Sigma}^{2}\bm{Q}\|_{F}$ $=\|\bm{Q}\left(\bm{\Delta}+a\bm{I}\right)^{2}-\left(\bm{\Delta}+a\bm{I}\right)^{2}\bm{Q}\|_{F}$
$=\left\|\bm{Q}\bm{\Delta}^{2}-\bm{\Delta}^{2}\bm{Q}+2a\left(\bm{Q}\bm{\Delta}-\bm{\Delta}\bm{Q}\right)\right\|_{F}$
where the second inequality follows from $\bm{Q}\bm{\Delta}-\bm{\Delta}\bm{Q}=\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}$ and $\|\bm{Q}\bm{\Delta}^{2}-\bm{\Delta}^{2}\bm{Q}\|_{F}=\|\bm{\Delta}(\bm{\Delta}\bm{Q}-\bm{Q}\bm{\Delta})+(\bm{\Delta}\bm{Q}-\bm{Q}\bm{\Delta})\bm{\Delta}\|_{F}\leq 2\|\bm{\Delta}\|\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\|_{F}$, and the last inequality uses $\|\bm{\Delta}\|\leq a/2$. ∎
B.3 Auxiliary Lemma and Proof for Proving 5
Lemma B.5.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\mathrm{sort}}\setminus\{\bm{0}\}$ and $i\in[r_{\sigma}]$ be arbitrary. Suppose that $|x-\sigma_{i}^{*}|\leq\min\left\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\right\},$ where $\delta_{1}$ and $\delta_{2}$ are defined in (87). Then we have
$\left|\varphi(x)-\varphi(\sigma_{i}^{*})\right|\leq\frac{\delta_{y}}{3}\text{ and }\ |\varphi^{\prime}(x)-\varphi^{\prime}(\sigma^{*}_{i})|\leq\frac{|\varphi^{\prime}(\sigma_{i}^{*})|}{2}.$ (133)
Proof.
According to the definition of $\delta_{\sigma}$ in (28), if $|x-\sigma_{i}^{*}|\leq\frac{\delta_{\sigma}}{3}$, then we have $\frac{2\sigma_{\min}^{*}}{3}\leq|x|\leq\frac{4\sigma_{\max}^{*}}{3}.$ Consequently, we obtain that
$|\varphi^{\prime}(x)|$ $=\Bigl|\frac{L}{\sqrt{\lambda}}x^{L-1}+\sqrt{\lambda}(2-L)x^{1-L}\Bigr|\leq\frac{L}{\sqrt{\lambda}}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}+(L-2)\sqrt{\lambda}\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{1-L},$
$|\varphi^{\prime\prime}(x)|$ $=\Bigl|\frac{L(L-1)}{\sqrt{\lambda}}x^{L-1}+\sqrt{\lambda}(2-L)(1-L)x^{-L}\Bigr|$
$\leq\frac{L(L-1)}{\sqrt{\lambda}}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}+\sqrt{\lambda}(2-L)(1-L)\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{-L}.$
Finally, by the mean value theorem, for all $x$ satisfying $|x-\sigma_{i}^{*}|\leq\min\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\}$, we obtain (133). ∎
B.3.1 Proof of (97) and (98)
Proof.
Using the form of $\bm{\Sigma}_{l}$ in (3.2.1) and (71) for $L=2$ (resp., (73) for $L\geq 3$), we have for each $l\in[L]$,